OpenStudy (anonymous):

can someone please check over my work I have the answer's I just don't understand the last part and I would love if someone could explain

OpenStudy (anonymous):

The braking distance needed to stop a car varies directly with the square of the car's speed. A car going 30 mph needs 40 feet of braking distance to come to a complete stop. How many feet of braking distance does a car need, under the same conditions, going 60 mph? d = ks2 d = 40 and s = 30 d = ks2 40 = k(30)2 40 = k(900) 900/900 gets canceled out d=2/45s2 that's what I can so far

OpenStudy (anonymous):

d=2/45s^2 d=2/24(6)^2 d=2/(3600) d=160 feet

OpenStudy (anonymous):

I just dont understand how it went from 2/45(3600) to 160 feet can someone please explain

ganeshie8 (ganeshie8):

looks correct

ganeshie8 (ganeshie8):

\(d = \frac{2}{45}s^2\)

OpenStudy (anonymous):

but can you please explain how it went from 2/45(3600) to 160 feet what did they divide by? I'm learning this in school so I just got a little confused on that part and I would love to understand

ganeshie8 (ganeshie8):

plugin s = 60

ganeshie8 (ganeshie8):

\(d = \frac{2}{45} \times 60^2\) \(d = \frac{2}{45} \times 3600\)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

yes I understand that part the only thing i'm not getting is the end where the answer is 160 feet because it went from 3600 to 160 how did they get that?

ganeshie8 (ganeshie8):

\(\large d = \frac{2}{\cancel{45}^5} \times \cancel{3600}^{400}\)

ganeshie8 (ganeshie8):

9 goes 5 times in 45, and 400 times in 3600

ganeshie8 (ganeshie8):

\(\large d = \frac{2}{\cancel{45}^\cancel{5}} \times \cancel{3600}^{\cancel{400}^{80}}\)

ganeshie8 (ganeshie8):

5 goes in 400, 80 times

ganeshie8 (ganeshie8):

leaving you with :- \(d = 2 \times 80\)

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

it does thank you(: so basically you just look for a number that can go into both of them and that number happened to be 9? i'm sorry I get very mixed up in math thank you for helping me

ganeshie8 (ganeshie8):

Exactly !

OpenStudy (anonymous):

okay awesome! it feels so much better finally understanding something(: thank you for helping me! I would give you a medal but I don't know how to

ganeshie8 (ganeshie8):

glad to hear that :) you're wlcme :D

ganeshie8 (ganeshie8):

you can click "Best Response" when you think somebody helped u

OpenStudy (anonymous):

okay I will do that(: thank you

ganeshie8 (ganeshie8):

np :)